The price of a share follows the geometric Brownian motion with parameters µ = 0.2 and σ = 0.18. Presently, the share’s price is £38. Consider a call option having one year until its expiration time and having a strike price of £40. The continuously compounded interest rate is 5%.

(a) What is the risk-neutral price C of this call option?

(b) Suppose now that you are the seller of this option. At time t = 0 you get £C from the buyer of the option, where C is the risk-neutral price of the option. You then have to design a hedging strategy which would allow you to meet your financial obligation in one year’s time. Your portfolio should consist of two investments: you are allowed to buy the underlying shares and to deposit money in the bank.

(i) The price of the share evolves according to a geometric Brownian motion. State the formulae you will need to compute the number of shares in the portfolio and the capital deposited in the bank at any time t, 0 6 t 6 1.


1
Expert's answer
2021-05-11T12:03:59-0400


a:

Risk-Neutral price of call option

"C=\frac{π\times c^{+}+(1-π)\times c^{-}}{1+r}"

where π is the probability, r is the risk-free rate

First, we need to determine the probability of up movement with help of the risk-neutral formula

"π=\frac{1+r-d}{u-d}"

where is the probability, r is the risk-free rate, u is the underlying price of the up move.

Substituting the values in the equation:

"π=\frac{1+5\%-0.18}{0.2-0.18} =\frac{0.87}{0.02}"

"\u03c0=43.5"



Substituting the values on the equation of risk-Neutral price of call option:

"c=\frac{43.5\times40+1-43.5\times38}{1+0.05}"

"c=\frac{125}{1.05}"

 "c=119.04"

Therefore, the risk-neutral price of the call option is $119.04.

(b)(i)

To compute the number of shares in a portfolio and the capital deposited in the bank at any time t, 0 ≤ t ≤ 1 in geometric Brownian motion, given equations are followed-

"\u2202V\u2202t+12\u03c32S2\u2202V\u2202S2+rS\u2202V\u2202S\u2212rV=0"

C(S,t), of this equation is equal to the cost of constructing an option from the specified stock. Here, r and σ are constant, this is:

"C(S,t)=SN(d1)\u2212Le\u2212r(T\u2212t)N(d2)"

L is the option's ‘strike price’

T is its time to maturity

N is the cumulative probability distribution function for a normal random variable.

Now-

"d1=log(S\/L)+(r+12\u03c32)(T\u2212t)\u03c3(T\u2212t)"

"d2=log(S\/L)+(r\u221212\u03c32)(T\u2212t)\u03c3(T\u2212t)=d1\u2212\u03c3(T\u2212t)"